# Signal attenuation enables scalable decentralized multi-agent reinforcement learning over networks

### v1

@@ III-B Signal Attenuation and Decentralization
Proof sketch. The main effort in the proof revolves around the case where $f^{i}=r^{i}$, since Assumption 2 renders the $f^{i}=c^{i}$ case trivial. Let $s=(s^{\mathcal{N}_{\kappa}(i)},s^{\mathcal{N}_{\kappa}^{-1}(i)}),s^{\prime}=(s^{\mathcal{N}_{\kappa}(i)},\bar{s}^{\mathcal{N}_{\kappa}^{-1}(i)})$, $a=(a^{\mathcal{N}_{\kappa}(i)},a^{\mathcal{N}_{\kappa}^{-1}(i)}),a^{\prime}=(a^{\mathcal{N}_{\kappa}(i)},\bar{a}^{\mathcal{N}_{\kappa}^{-1}(i)})$. In the $f^{i}=r^{i}$ case, we begin by bounding the absolute value of the worst-case difference, $\sup|r^{i}(s,a)-r^{i}_{\kappa}(s^{\prime},a^{\prime})|$, between $r^{i}$ defined in (5) and a local version of the same reward, $r^{i}_{\kappa}$, depending only on information from $\mathcal{N}_{\kappa}(i)$ and defined by
@@
$r^{i}_{\kappa}(s_{t},a_{t})=\frac{h^{\tau}_{ii}(s_{t})a_{t}^{i}}{\sigma^{2}(s_{t},a_{t})+\sum_{j\in\mathcal{N}_{\kappa}(i)}c_{ji}(s_{t})\left(h^{d}_{ji}(s_{t})+h^{\tau}_{ji}(s_{t})\right)a_{t}^{j}}.$ (14)
This bound is obtained by leveraging Assumption 1, Assumption 3, and path loss properties arising from the radar range equations (2)-(3). With this bound in hand, we next bound $|Q^{r^{i}}(s,a)-Q^{r^{i}}(s^{\prime},a^{\prime})|$ by a product of the bound on $\sup|r^{i}(s,a)-r^{i}_{\kappa}(s^{\prime},a^{\prime})|$ and a total variation distance term between the state occupancy measures induced by $\pi_{\theta}$ from the different initializations $s$ and $s^{\prime}$. To finish, we bound this total variation distance using Assumptions 1 and 4. $\square$
@@
We now proceed to the main result of this section, which establishes bounds on the accuracy of gradient estimators constructed using only local neighborhood information. We first define local $Q$ function approximations. Fix $i\in\mathcal{N}$, and let $w_{i}:\mathcal{S}^{\mathcal{N}^{-1}_{\kappa}(i)}\times\mathcal{A}^{\mathcal{N}^{-1}_{\kappa}(i)}\rightarrow[0,1]$ be an arbitrary weighting function satisfying $\sum_{\bar{s}\in\mathcal{S}^{\mathcal{N}^{-1}_{\kappa}(i)},\bar{a}\in\mathcal{A}^{\mathcal{N}^{-1}_{\kappa}(i)}}w_{i}(\bar{s},\bar{a})=1$. Let $\widetilde{Q}_{\theta}^{f^{i}}(s^{\mathcal{N}_{\kappa}(i)},a^{\mathcal{N}_{\kappa}(i)})$ denote agent $i$’s local approximation of $Q^{f^{i}}(s,a)$, for $f^{i}\in\{r^{i},c^{i}\}$, defined by
@@
$\sum_{\bar{s}\in\mathcal{S}^{\mathcal{N}^{-1}_{\kappa}(i)},\bar{a}\in\mathcal{A}^{\mathcal{N}^{-1}_{\kappa}(i)}}Q_{\theta}^{f^{i}}\left((s^{\mathcal{N}_{\kappa}(i)},\bar{s}),(a^{\mathcal{N}_{\kappa}(i)},\bar{a})\right)w_{i}(\bar{s},\bar{a}).$ (15)
For any such weighting function $w_{i}$, the following approximation result is implied by Theorem 1.
@@
Let the conditions of Theorem 1 and Assumption 5 hold. Fix $i,j\in\mathcal{N},f^{i}\in\{r^{i},c^{i}\}$, and define $f(s,a)=\sum_{i\in\mathcal{N}}f^{i}(s^{i},a^{i})$. We have the following:
@@
    $|\widetilde{Q}_{\theta}^{f^{i}}(s^{\mathcal{N}_{\kappa}(i)},a^{\mathcal{N}_{\kappa}(i)})-Q^{f^{i}}(s,a)|\leq\frac{M|\mathcal{N}_{\kappa}^{-1}(i)|}{g^{2}(\kappa,R)}$,
@@
    $\left\lVert\widehat{h_{f^{j}}^{i}}(\theta)-\nabla_{\theta^{i}}J_{f^{j}}(\theta)\right\rVert\leq\frac{ML^{i}|\mathcal{N}_{\kappa}^{-1}(i)|}{g^{2}(\kappa,R)}$, where $\widehat{h^{i}_{f^{j}}}(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\widetilde{Q}_{\theta}^{f^{j}}(s^{\mathcal{N}_{\kappa}(j)},a^{\mathcal{N}_{\kappa}(j)})\nabla_{\theta^{i}}\log\pi^{i}_{\theta^{i}}(a^{i}|s^{i})\right]$;
@@
    $\left\lVert\widehat{h^{i}_{f}}(\theta)-\nabla_{\theta^{i}}J_{f}(\theta)\right\rVert\leq\frac{MnL^{i}|\mathcal{N}_{\kappa}^{-1}(i)|}{g^{2}(\kappa,R)},$ where $\widehat{h^{i}_{f}}(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{j\in\mathcal{N}_{\kappa}(i)}\widetilde{Q}_{\theta}^{f^{j}}(s^{\mathcal{N}_{\kappa}(j)},a^{\mathcal{N}_{\kappa}(j)})\nabla_{\theta^{i}}\log\pi^{i}_{\theta^{i}}(a^{i}|s^{i})\right]$.
@@
    If, in addition, Assumption 6 holds, then we have $\left\lVert\sum_{j\in\mathcal{N}_{\kappa}(i)}\lambda^{j}\widehat{h^{i}_{f^{j}}}(\theta)-\sum_{l\in\mathcal{N}}\lambda^{l}\nabla_{\theta^{i}}J_{f^{l}}(\theta)\right\rVert$

### v2

@@ III-B Signal Attenuation and Decentralization
Proof.
Assumption 2 renders the $f^{i}=c^{i}$ case trivial, so we give the proof only for the $f^{i}=r^{i}$ case. Fix $\theta\in\Theta,i\in\mathcal{N},$ and $\kappa\in\mathbb{N}^{+}$. Let $s^{\mathcal{N}_{\kappa}(i)}\in\mathcal{S}^{\mathcal{N}_{\kappa}(i)}$, $a^{\mathcal{N}_{\kappa}(i)}\in\mathcal{A}^{\mathcal{N}_{\kappa}(i)}$, $s^{\mathcal{N}^{-1}_{\kappa}(i)},\bar{s}^{\mathcal{N}^{-1}_{\kappa}(i)}\in\mathcal{S}^{\mathcal{N}^{-1}_{\kappa}(i)}$, and $a^{\mathcal{N}^{-1}_{\kappa}(i)},\bar{a}^{\mathcal{N}^{-1}_{\kappa}(i)}\in\mathcal{A}^{\mathcal{N}^{-1}_{\kappa}(i)}$. Define $s=(s^{\mathcal{N}_{\kappa}(i)},s^{\mathcal{N}_{\kappa}^{-1}(i)}),\bar{s}=(s^{\mathcal{N}_{\kappa}(i)},\bar{s}^{\mathcal{N}_{\kappa}^{-1}(i)}),a=(a^{\mathcal{N}_{\kappa}(i)},a^{\mathcal{N}_{\kappa}^{-1}(i)})$, and $\bar{a}=(a^{\mathcal{N}_{\kappa}(i)},\bar{a}^{\mathcal{N}_{\kappa}^{-1}(i)})$.
By the definition of $Q^{r^{i}}_{\theta}$, we have
$\left\lvert Q^{r^{i}}_{\theta}(s,a)-Q^{r^{i}}_{\theta}(\bar{s},\bar{a})\right\rvert$ (16)
$=\big{|}\mathbb{E}_{\pi_{\theta}}\big{[}\sum_{t=0}^{\infty}r^{i}(s_{t},a_{t})-J_{r^{i}}(\theta)\ |\ s_{0}=s,a_{0}=a\big{]}$
$\qquad-\mathbb{E}_{\pi_{\theta}}\big{[}\sum_{t=0}^{\infty}r^{i}(s_{t},a_{t})-J_{r^{i}}(\theta)\ |\ s_{0}=\bar{s},a_{0}=\bar{a}\big{]}\big{|}$ (17)
$=\big{|}\mathbb{E}_{\pi_{\theta}}\big{[}\sum_{t=0}^{\infty}r^{i}(s_{t},a_{t})\ |\ s_{0}=s,a_{0}=a\big{]}$
$\qquad-\mathbb{E}_{\pi_{\theta}}\big{[}\sum_{t=0}^{\infty}r^{i}(s_{t},a_{t})\ |\ s_{0}=\bar{s},a_{0}=\bar{a}\big{]}\big{|},$ (18)
where the last equality holds by the fact that $\mathbb{E}_{\pi_{\theta}}\left[J_{r^{i}}(\theta)\ |\ s_{0}=s,a_{0}=a\right]=\mathbb{E}_{\pi_{\theta}}\left[J_{r^{i}}(\theta)\ |\ s_{0}=\bar{s},a_{0}=\bar{a}\right]$, since Assumption 4 implies ergodicity of the Markov chain over $\mathcal{S}$ induced by $\pi_{\theta}$. Recalling (6) and continuing from (18), we have
$\left\lvert Q^{r^{i}}_{\theta}(s,a)-Q^{r^{i}}_{\theta}(\bar{s},\bar{s})\right\rvert$ (19)
$=\Big{|}\sum_{t=0}^{\infty}\Big{[}\mathbb{E}_{\pi_{\theta}}\big{[}r^{i}(s_{t},a_{t})-r^{i}_{\kappa}(s_{t},a_{t})+r^{i}_{\kappa}(s_{t},a_{t})\ |\ s_{0}=s,a_{0}=a\big{]}$
$\qquad-\mathbb{E}_{\pi_{\theta}}\big{[}r^{i}(s_{t},a_{t})\ |\ s_{0}=\bar{s},a_{0}=\bar{a}\big{]}\Big{]}\Big{|}$ (20)
$=\Big{|}\sum_{t=0}^{\infty}\Big{[}\mathbb{E}_{\pi_{\theta}}\big{[}r^{i}(s_{t},a_{t})-r^{i}_{\kappa}(s_{t},a_{t})\ |\ s_{0}=s,a_{0}=a\big{]}$
$\qquad-\mathbb{E}_{\pi_{\theta}}\big{[}r^{i}(s_{t},a_{t})-r^{i}_{\kappa}(s_{t},a_{t})\ |\ s_{0}=\bar{s},a_{0}=\bar{a}\big{]}\Big{]}\Big{|},$ (21)
where (21) follows by the independence of $a$ of the transition dynamics $p$ defined in Section II-B and the fact that $r^{i}_{\kappa}(s,a)$ is independent of the value of $s^{\mathcal{N}_{\kappa}^{-1}(i)},a^{\mathcal{N}_{\kappa}^{-1}(i)}$ by definition.
Now notice that
$\sum_{t=0}^{\infty}\mathbb{E}_{\pi_{\theta}}\big{[}r^{i}(s_{t},a_{t})-r^{i}_{\kappa}(s_{t},a_{t})\ |\ s_{0}=s,a_{0}=a\big{]}$ (22)
$=\sum_{t=0}^{\infty}\sum_{s^{\prime}\in\mathcal{S}}\sum_{a^{\prime}\in\mathcal{A}}G^{i}_{\kappa}(s^{\prime},a^{\prime})\pi_{\theta}(a^{\prime}|s^{\prime})d_{\pi_{\theta}}^{t}(s^{\prime}|s_{0}=s,a_{0}=a),$ (23)
where $d_{\pi_{\theta}}^{t}$ denotes the $t$-step state distribution of Assumption 4 and $G^{i}_{\kappa}(s^{\prime},a^{\prime})=r^{i}(s^{\prime},a^{\prime})-r^{i}_{\kappa}(s^{\prime},a^{\prime})$, and that an analogous expression holds for $s_{0}=\bar{s},a_{0}=\bar{a}$. Continuing from (21) and (23), we have
$\left\lvert Q^{r^{i}}_{\theta}(s,a)-Q^{r^{i}}_{\theta}(\bar{s},\bar{s})\right\rvert$ (24)
$=\Big{|}\sum_{t=0}^{\infty}\sum_{s^{\prime}\in\mathcal{S}}\sum_{a^{\prime}\in\mathcal{A}}G_{\kappa}^{i}(s^{\prime},a^{\prime})\pi_{\theta}(a^{\prime}|s^{\prime})\Big{(}d_{\pi_{\theta}}^{t}(s^{\prime}|s_{0}=s,a_{0}=a)$
$\hskip 113.81102pt-d_{\pi_{\theta}}^{t}(s^{\prime}|s_{0}=\bar{s},a_{0}=\bar{a})\Big{)}\Big{|}$ (25)
$\stackrel{{\scriptstyle\textnormal{(a)}}}{{\mathstrut{\leq}}}\sum_{t=0}^{\infty}\sum_{s^{\prime}\in\mathcal{S}}\sum_{a^{\prime}\in\mathcal{A}}\left\lvert G^{i}_{\kappa}(s^{\prime},a^{\prime})\right\rvert\cdot\big{|}d_{\pi_{\theta}}^{t}(s^{\prime}|s_{0}=s,a_{0}=a)$
$\hskip 113.81102pt-d_{\pi_{\theta}}^{t}(s^{\prime}|s_{0}=\bar{s},a_{0}=\bar{a})\big{|}$ (26)
$\stackrel{{\scriptstyle\textnormal{(b)}}}{{\mathstrut{\leq}}}\sum_{t=0}^{\infty}\sum_{s^{\prime}\in\mathcal{S}}\sum_{a^{\prime}\in\mathcal{A}}\left\lvert G^{i}_{\kappa}(s^{\prime},a^{\prime})\right\rvert\Big{[}\left\lvert d_{\pi_{\theta}}^{t}(s^{\prime}|s_{0}=s,a_{0}=a)-d_{\pi_{\theta}}(s^{\prime})\right\rvert$
$\hskip 56.9055pt+\left\lvert d_{\pi_{\theta}}^{t}(s^{\prime}|s_{0}=\bar{s},a_{0}=\bar{a})-d_{\pi_{\theta}}(s^{\prime})\right\rvert\Big{]}$ (27)
$\stackrel{{\scriptstyle\textnormal{(c)}}}{{\mathstrut{\leq}}}\sup_{s^{\prime},a^{\prime}}\left\lvert G^{i}_{\kappa}(s^{\prime},a^{\prime})\right\rvert\sum_{t=0}^{\infty}\Big{[}d_{TV}(d_{\pi_{\theta}}^{t}(\cdot|s_{0}=s,a_{0}=a)\ ||\ d_{\pi_{\theta}}(\cdot))$
$\hskip 56.9055pt+d_{TV}(d_{\pi_{\theta}}^{t}(\cdot|s_{0}=\bar{s},a_{0}=\bar{a})\ ||\ d_{\pi_{\theta}}(\cdot))\Big{]}$ (28)
$\stackrel{{\scriptstyle\textnormal{(d)}}}{{\mathstrut{\leq}}}\sup_{s^{\prime},a^{\prime}}\left\lvert G^{i}_{\kappa}(s^{\prime},a^{\prime})\right\rvert\sum_{t=0}^{\infty}2m\rho^{t}=\sup_{s^{\prime},a^{\prime}}\left\lvert G^{i}_{\kappa}(s^{\prime},a^{\prime})\right\rvert\frac{2m}{1-\rho},$ (29)
where inequality (a) follows by the triangle and Cauchy-Schwarz inequalities and the fact that $0\leq\pi_{\theta}(a^{\prime}|s^{\prime})\leq 1$, inequality (b) follows by adding and subtracting $d_{\theta}(s^{\prime})$ and applying the triangle inequality, (c) follows by taking suprema and the definition of total variation distance, and (d) follows by Assumption 4.
All that remains is to bound $\sup_{s^{\prime},a^{\prime}}\left\lvert G^{i}_{\kappa}(s^{\prime},a^{\prime})\right\rvert$. Fix $s\in\mathcal{S},a\in\mathcal{A}$. Recall that
$r^{i}(s,a)=\frac{h^{\tau}_{ii}(s)a^{i}}{\sigma^{2}(s,a)+\alpha},\quad r^{i}_{\kappa}(s,a)=\frac{h^{\tau}_{ii}(s)a^{i}}{(\sigma_{\kappa}(s,a))^{2}+\beta},$ (30)
where
$\alpha$ $=\sum_{j\in\mathcal{N}\setminus\{i\}}c_{ji}(s)\left[h^{d}_{ji}(s)+h^{\tau}_{ji}(s)\right]a^{j},$ (31)
$\beta$ $=\sum_{j\in\mathcal{N}_{\kappa}(i)\setminus\{i\}}c_{ji}(s)\left[h^{d}_{ji}(s)+h^{\tau}_{ji}(s)\right]a^{j}.$ (32)
In light of this, we can write
@@
$\left\lvert G^{i}_{\kappa}(s,a)\right\rvert$ $=\left\lvert\frac{h^{\tau}_{ii}(s)a^{i}(\alpha-\beta)}{(\sigma^{2}(s,a)+\alpha)((\sigma^{i}_{\kappa}(s,a))^{2}+\beta)}\right\rvert$ (33)
=$\stackrel{{\scriptstyle\textnormal{(e)}}}{{\mathstrut{=}}}\frac{h^{\tau}_{ii}(s)a^{i}\left\lvert\alpha-\beta\right\rvert}{\left\lvert\sigma^{2}(s,a)+\alpha\right\rvert\left\lvert(\sigma^{i}_{\kappa}(s,a))^{2}+\beta\right\rvert}$ (34)
$\stackrel{{\scriptstyle\textnormal{(f)}}}{{\mathstrut{\leq}}}\frac{h^{\tau}_{ii}(s)a^{i}\left\lvert\alpha-\beta\right\rvert}{\underset{\bar{}}{\sigma}^{4}},$ (35)
where (a) follows by nonnegativity of $h^{\tau}_{ii}(s)$ and $a^{i}$, and (b) follows by nonnegativity of $\alpha$ and $\beta$ and Assumption 3. By (2) and Assumption 3,
$h^{\tau}_{ii}(s)a^{i}=\frac{G_{t}G_{r}\sigma^{RCS}_{ii}(s)\lambda^{2}a^{i}}{(4\pi)^{3}R_{i}^{4}(s)}\leq\frac{G_{t}G_{r}\bar{\sigma}^{RCS}\lambda^{2}a^{max}}{(4\pi)^{3}},$ (36)
whence it follows that
$\left\lvert G^{i}_{\kappa}(s,a)\right\rvert\leq\frac{G_{t}G_{r}\bar{\sigma}^{RCS}\lambda^{2}a^{max}}{(4\pi)^{3}\underset{\bar{}}{\sigma}^{4}}\left\lvert\alpha-\beta\right\rvert.$ (37)
It remains to bound $\left\lvert\alpha-\beta\right\rvert$. Notice that
$\alpha-\beta$ $=\sum_{j\in\mathcal{N}_{\kappa}^{-1}(i)}c_{ji}(s)\left[h^{d}_{ji}(s)+h^{\tau}_{ji}(s)\right]a^{j}$ (38)
$\leq\bar{c}a^{max}\sum_{j\in\mathcal{N}_{\kappa}^{-1}(i)}\left[h^{d}_{ji}(s)+h^{\tau}_{ji}(s)\right]$ (39)
so we need only bound $h^{d}_{ji}(s)$ and $h^{\tau}_{ji}(s)$. For $h^{d}_{ji}(s)$, notice that, for $j\in\mathcal{N}_{\kappa}^{-1}(i)$, by (3) and Assumption 1 we have
$h^{d}_{ji}(s)=\frac{G_{t}^{\prime}G_{r}^{\prime}\lambda^{2}}{(4\pi)^{2}(d_{ij}(s))^{2}}\leq\frac{G_{t}^{\prime}G_{r}^{\prime}\lambda^{2}}{(4\pi)^{2}g^{2}(\kappa,R)}.$ (40)
For $h^{\tau}_{ji}(s)$, we consider four possible cases that may arise depending on the physical configuration of the radar network. Fix $j\in\mathcal{N}_{\kappa}^{-1}(i)$.
Case 1: $R_{j}(s)\leq d_{ij}(s)\leq R_{i}(s)$. In this case, we know $R_{i}(s)R_{j}(s)\geq d_{ij}(s)$, whence
$h^{\tau}_{ji}(s)=\frac{G_{t}G_{r}\sigma^{RCS}(s)\lambda^{2}}{(4\pi)^{3}R^{2}_{i}(s)R^{2}_{j}(s)}\leq\frac{G_{t}G_{r}\bar{\sigma}^{RCS}\lambda^{2}}{(4\pi)^{3}g^{2}(\kappa,R)},$ (41)
where the inequality holds by Assumptions 1 and 3.
Case 2: $R_{i}(s)\leq d_{ij}(s)\leq R_{j}(s)$. This case follows by reasoning identical to Case 1.
Case 3: $d_{ij}(s)\leq\min\{R_{i}(s),R_{j}(s)\}$. In this case, $R_{i}(s)R_{j}(s)\geq d^{2}_{ij}(s)$, whence
$h^{\tau}_{ij}(s)\leq\frac{G_{t}G_{r}\bar{\sigma}^{R}CS\lambda^{2}}{(4\pi)^{3}g^{4}(\kappa,R)}\stackrel{{\scriptstyle\textnormal{(g)}}}{{\mathstrut{\leq}}}\frac{G_{t}G_{r}\bar{\sigma}^{R}CS\lambda^{2}}{(4\pi)^{3}g^{2}(\kappa,R)},$ (42)
where (a) holds since $g(\kappa,R)\geq 1$ by Assumption 1.
Case 4: $d_{ij}(s)\geq\max\{R_{i}(s),R_{j}(s)\}$. The triangle inequality gives $R_{i}(s)+R_{j}(s)\geq d_{ij}(s)$, whence $R_{i}(s)\geq g(\kappa,R)-R_{j}(s)$. If $R_{j}\geq\tfrac{1}{2}g(\kappa,R)$, then
$h^{\tau}_{ij}(s)$ $=\frac{G_{t}G_{r}\sigma^{RCS}(s)\lambda^{2}}{(4\pi)^{3}R^{2}_{i}(s)R^{2}_{j}(s)}\leq\frac{G_{t}G_{r}\sigma^{RCS}(s)\lambda^{2}}{(4\pi)^{3}R^{2}_{j}(s)}$ (43)
$\leq\frac{G_{t}G_{r}\bar{\sigma}^{RCS}\lambda^{2}}{(4\pi)^{3}\tfrac{1}{4}g^{2}(\kappa,R)}.$ (44)
On the other hand, if $R_{j}(s)<\tfrac{1}{2}g(\kappa,R)$, then $R_{i}(s)\geq g(\kappa,R)-R_{j}(s)>\tfrac{1}{2}g(\kappa,R)$. Thus, by the same argument as above,
$h^{\tau}_{ji}(s)\leq\frac{G_{t}G_{r}\bar{\sigma}^{RCS}\lambda^{2}}{(4\pi)^{3}\tfrac{1}{4}g^{2}(\kappa,R)}.$ (45)
Taking the maximum over all four cases, we obtain
$h^{\tau}_{ji}(s)\leq\frac{G_{t}G_{r}\bar{\sigma}^{RCS}\lambda^{2}}{4^{2}\pi^{3}g^{2}(\kappa,R)}.$ (46)
Combining (39), (40), and (46), we have
$\left\lvert\alpha-\beta\right\rvert\leq\bar{c}a^{max}\left\lvert\mathcal{N}_{\kappa}^{-1}(i)\right\rvert\left[\frac{G_{t}^{\prime}G_{r}^{\prime}\lambda^{2}}{(4\pi)^{2}g^{2}(\kappa,R)}+\frac{G_{t}G_{r}\bar{\sigma}^{RCS}\lambda^{2}}{4^{2}\pi^{3}g^{2}(\kappa,R)}\right],$ (47)
which, combined with (37), gives us
$\left\lvert G^{i}_{\kappa}(s,a)\right\rvert$ $\leq\frac{G_{t}G_{r}\bar{\sigma}^{RCS}\lambda^{2}(a^{max})^{2}\bar{c}}{(4\pi)^{3}}\cdot$
$\hskip 14.22636pt\left[\frac{G_{t}^{\prime}G_{r}^{\prime}\lambda^{2}}{(4\pi)^{2}}+\frac{G_{t}G_{r}\bar{\sigma}^{RCS}\lambda^{2}}{4^{2}\pi^{3}}\right]\frac{\left\lvert\mathcal{N}_{\kappa}^{-1}(i)\right\rvert}{g^{2}(\kappa,R)}.$ (48)
Finally, combining (29) with (48), we have
$\left\lvert Q^{r^{i}}_{\theta}(s,a)-Q^{r^{i}}_{\theta}(\bar{s},\bar{s})\right\rvert\leq\frac{M\left\lvert\mathcal{N}_{\kappa}^{-1}(i)\right\rvert}{g^{2}(\kappa,R)},$ (49)
where
$M=\frac{2mG_{t}G_{r}\bar{\sigma}^{RCS}\lambda^{4}(a^{max})^{2}\bar{c}}{(1-\rho)(4\pi)^{3}}\left[\frac{G_{t}^{\prime}G_{r}^{\prime}}{(4\pi)^{2}}+\frac{G_{t}G_{r}\bar{\sigma}^{RCS}}{4^{2}\pi^{3}}\right].$ (50)
@@
We now proceed to the main result of this section, which establishes bounds on the accuracy of gradient estimators constructed using only local neighborhood information. We first define local $Q$ function approximations. Fix $i\in\mathcal{N}$, and let $w^{i}:\mathcal{S}^{\mathcal{N}^{-1}_{\kappa}(i)}\times\mathcal{A}^{\mathcal{N}^{-1}_{\kappa}(i)}\rightarrow[0,1]$ be an arbitrary weighting function satisfying $\sum_{\bar{s}\in\mathcal{S}^{\mathcal{N}^{-1}_{\kappa}(i)},\bar{a}\in\mathcal{A}^{\mathcal{N}^{-1}_{\kappa}(i)}}w^{i}(\bar{s},\bar{a})=1$. Let $\widetilde{Q}_{\theta}^{f^{i}}(s^{\mathcal{N}_{\kappa}(i)},a^{\mathcal{N}_{\kappa}(i)})$ denote agent $i$’s local approximation of $Q^{f^{i}}(s,a)$, for $f^{i}\in\{r^{i},c^{i}\}$, defined by
@@
$\sum_{\bar{s}\in\mathcal{S}^{\mathcal{N}^{-1}_{\kappa}(i)},\bar{a}\in\mathcal{A}^{\mathcal{N}^{-1}_{\kappa}(i)}}Q_{\theta}^{f^{i}}\left((s^{\mathcal{N}_{\kappa}(i)},\bar{s}),(a^{\mathcal{N}_{\kappa}(i)},\bar{a})\right)w^{i}(\bar{s},\bar{a}).$ (51)
For any such weighting function $w^{i}$, the following approximation result is implied by Theorem 1.
@@
Let the conditions of Theorem 1 and Assumption 5 hold. Fix $i,j\in\mathcal{N},f^{i}\in\{r^{i},c^{i}\}$, define $f(s,a)=\sum_{i\in\mathcal{N}}f^{i}(s^{i},a^{i})$, and let $M$ be as in Theorem 1. We have:
@@
    $|\widetilde{Q}_{\theta}^{f^{i}}(s^{\mathcal{N}_{\kappa}(i)},a^{\mathcal{N}_{\kappa}(i)})-Q_{\theta}^{f^{i}}(s,a)|\leq\frac{M|\mathcal{N}_{\kappa}^{-1}(i)|}{g^{2}(\kappa,R)}$,
@@
    $\left\lVert\widehat{h_{f^{j}}^{i}}(\theta)-\nabla_{\theta^{i}}J_{f^{j}}(\theta)\right\rVert\leq\frac{ML^{i}|\mathcal{N}_{\kappa}^{-1}(j)|}{g^{2}(\kappa,R)}$, where $\widehat{h^{i}_{f^{j}}}(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\widetilde{Q}_{\theta}^{f^{j}}(s^{\mathcal{N}_{\kappa}(j)},a^{\mathcal{N}_{\kappa}(j)})\nabla_{\theta^{i}}\log\pi^{i}_{\theta^{i}}(a^{i}|s^{i})\right]$;
@@
    If Assumption 6 also holds, then $\left\lVert\widehat{h^{i}_{f}}(\theta)-\nabla_{\theta^{i}}J_{f}(\theta)\right\rVert\leq\frac{M\bar{n}L^{i}|\mathcal{N}_{\kappa}^{-1}(i)|}{g^{2}(\kappa,R)}+\left\lvert\mathcal{N}_{\kappa}^{-1}(i)\right\rvert\varepsilon_{\kappa},$ where $\widehat{h^{i}_{f}}(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{j\in\mathcal{N}_{\kappa}(i)}\widetilde{Q}_{\theta}^{f^{j}}(s^{\mathcal{N}_{\kappa}(j)},a^{\mathcal{N}_{\kappa}(j)})\nabla_{\theta^{i}}\log\pi^{i}_{\theta^{i}}(a^{i}|s^{i})\right]$ and $\bar{n}=\max_{j\in\mathcal{N}}\left\lvert\mathcal{N}_{\kappa}^{-1}(j)\right\rvert$.
@@
    If Assumption 6 also holds, then $\left\lVert\sum_{j\in\mathcal{N}_{\kappa}(i)}\eta^{j}\widehat{h^{i}_{f^{j}}}(\theta)-\sum_{l\in\mathcal{N}}\eta^{l}\nabla_{\theta^{i}}J_{f^{l}}(\theta)\right\rVert$